OpenStudy (anonymous):

THE LENGTH OF arc EF is 5pi. what is the radius?

OpenStudy (anonymous):



OpenStudy (anonymous):

The circumference = (12)(5)(pi) = (60)(pi) because (30)(12) = 360. (60)(pi) = (2)(r)(pi) so, r = 30

OpenStudy (anonymous):

where did you get 12?

OpenStudy (anonymous):

30 degrees is 1/12 of the whole way around the circle. So, if you multiply the arc length by 12, you will get the circumference, the arc length of the whole circle. 30 = 360/12

OpenStudy (anonymous):

And then the circumference is (2)(pi)(r)

OpenStudy (anonymous):

oh! i get it!

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @Amenah